OpenStudy (anonymous):

The center of a circle is at (-10, 6) and it has a radius of 4. What is the equation of the circle? (Points : 4) (x -10)^2 + (y + 6)^2 = 2 (x -10)^2 + (y + 6)^2 = 16 (x + 10)^2 + (y - 6)^2 = 2 (x + 10)^2 + (y - 6)^2 = 16

OpenStudy (sepeario):

are you aware of how to find the equation or a circle?

OpenStudy (anonymous):

using the radius right?

Nnesha (nnesha):

\[\huge\rm (x-h)^2 + (y-k)^2 = r^2\] standard form equation of circle where (h,k) is the center and r is radius

OpenStudy (anonymous):

Oh yeah that does look really familiar

Nnesha (nnesha):

okay so replace (h,k) points by (-10,6) r=4

OpenStudy (solomonzelman):

An equation of the circle with center \(\large\color{black}{ \left(\color{red}{\rm h},\color{blue}{\rm k}\right) }\) and radius \(\large\color{green}{ \rm r }\) is: \(\large\color{black}{ (x-\color{red}{\rm h})^2+(y-\color{blue}{\rm k})^2=\color{green}{\rm r}^2 }\)

OpenStudy (solomonzelman):

took a little long to do the colors

OpenStudy (anonymous):

so its( x- 10^2 )+ (y- 6)^2 = 4

OpenStudy (solomonzelman):

^2 is outside the parenthesis

Nnesha (nnesha):

r=4 r^2 = ?

OpenStudy (anonymous):

oops it would be after that whats next?

OpenStudy (solomonzelman):

^2 next to the 4 on the right side.

OpenStudy (anonymous):

@Nnesha 16

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

Then?

OpenStudy (solomonzelman):

re-write it correctly please. with all corrections

OpenStudy (anonymous):

(x-10)^2 + (y-6)^2 = 16

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

good job, bye

OpenStudy (anonymous):

wait i still need help figuring it out

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @cookiimonster27 (x-10)^2 + (y-6)^2 = 16 \(\color{blue}{\text{End of Quote}}\) nope there is a negative sign in the parentheses and h is also negative so...

OpenStudy (anonymous):

itd be (x-(-10))^2 + (y-6)^2=16 @Nnesha which would make the 10 positive so its D

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

can you help with one last ine please

Nnesha (nnesha):

hmm i'll try

OpenStudy (anonymous):

The equation of a parabola is shown. y= 1/10x^2 What are the coordinates of the focus? (Points : 4) (0, 2.5) (0, -2.5) (0, 0.8) (0, -0.4) I think its D

Nnesha (nnesha):

it's horizontal or vertical ?

OpenStudy (anonymous):

 It goes like that

Nnesha (nnesha):

okay so it's ? vertical or horizontal ? :-)

Nnesha (nnesha):

remember the quadratic equation y =Ax^2+Bx+C ^^^ x is squared so

OpenStudy (anonymous):

I think vertical

Nnesha (nnesha):

yes right x is squared so it's vertical

OpenStudy (anonymous):

Yes x is squared .what do you think it is

Nnesha (nnesha):

so focus should be \[\huge\rm (0 , \pm P)\] do you know wat is p?

OpenStudy (anonymous):

I thhnk its 0.04

Nnesha (nnesha):

okay standard form is  4p should be with y so that's why first solve for x^2